hey
ok. the bases of a trapezoid are 3ft and 4ft. the area is 42 sq ft. i need help finding the height.
thanks

Answer by checkley75(3666)

H*(B1+B2)/2=AREA
H*(3+4)/2=42
H*7/2=42 NOW CROSS MULTIPLY
7H=42*2
7H=84
H=84/7
H=12 ANSWER
PROOF
12*(3+4)/2=42
12*7/2=42
84/2=42
42=42